OpenStudy (yanasidlinskiy):

Solve the equation. Show your work. 1/a = 6/18

OpenStudy (perl):

cross multiply 1*18 = a*6 divide both sides a = 3

OpenStudy (yanasidlinskiy):

thanx for your help:)